This is the question: Write the equation 3x-4y=20 in slope intercept form.
Help, thank you

Answer by usyim88hk(157)

When it says write it in slope intercept form, it means to solve for y
3x-4y = 20 (switch 20 with 4y)
3x-20 = 4y (divide both sides by 4)
y = (3/4)x - 5 (Answer!!!)
This is the slope intercept form